# Indecomposable involutive set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

### v2

@@ 3.1. Constructing the solutions
Let $0\neq f\in\operatorname{Fix}(\mathcal{G})$. Then ${}^{f}\mathcal{B}_{x}=\lambda_{f}(\mathcal{B}_{x})\neq\mathcal{B}_{x}$ as $f\not\in\mathcal{A}$. Therefore, $\operatorname{Fix}(\mathcal{G})$ acts transitively on the system of blocks by conjugation. Furthermore, by 3.6, $\operatorname{Fix}(\mathcal{G})$ acts transitively on every single block by right-multiplication. Therefore, each $y\in X$ is of the form $y={}^{f}x\circ f^{\prime}$ for some $f,f^{\prime}\in\operatorname{Fix}(\mathcal{G})$, which can be rewritten as $y=f\circ x\circ f^{\prime}\circ\overline{f}$. ∎
From 3.7 it therefore follows that we can coordinatize the elements of $X$ in the following way: by 3.6 we can choose an element $\alpha_{0,0}\in X\cap\mathcal{A}$ and some $0\neq f\in\operatorname{Fix}(\mathcal{G})$. For $(a,x)\in\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, we then set $\alpha_{a,x}=\left({}^{\overline{f}^{\circ a}}\alpha_{0,0}\right)\circ f^{\circ x}$. By 3.7, this gives a unique coordinatization of the elements in $X$.
@@
where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. Vice versa, this construction always results in an indecomposable, irretractable cycle set with $\mathcal{G}(X)$ a $p$-group.
@@
A comparison shows that in order for $X$ to satisfy Eq. C1, $\Phi(b-a)=\Phi(a-b)$ must hold for all $a,b\in\mathbb{Z}_{p}$ which amounts to saying that $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. By the same calculation, one sees that this obstruction to $\Phi$ is sufficient for $X$ to satisfy Eq. C1. By construction, all maps $\sigma_{(a,x)}$ are bijective. Furthermore, the square map
@@
Note that, using the explicit form in 3.8, we are considering the set $X\cap\mathcal{A}=\{(a,0)\mid a\in\mathbb{Z}_{p}\}$. As $\Phi$ is non-constant, we easily see that these elements generate the whole cycle set $X$. ∎
@@ 3.2. Getting rid of redundancy and determining automorphisms
Let $\mathcal{F}_{p}$ be the set of all non-constant functions $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ with the property that $\Phi(A)=\Phi(-A)$ for all $A\in\mathbb{Z}_{p}$. $\mathcal{F}_{p}$ is acted upon by $\mathbb{Z}_{p}^{\ast}$ via $({}^{\alpha}\Phi)(A)=\alpha^{-1}\Phi(\alpha A)$. From now on, let $\mathcal{R}_{p}$ be a fixed system of representatives for this action.
Recall from 3.8 that the cycle sets in the considered case are described as $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ with the operation
@@
By 3.2, $\mathcal{G}(X)^{\circ}$ is non-abelian and, by 1.7, has a unique non-trivial system of imprimitivity that consists of the blocks $\{a\}\times\mathbb{Z}_{p}$. Assume that $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ comes with two cycle set operations $\cdot,\cdot^{\prime}$ that are given by the parameters $\Phi,\Phi^{\prime}\in\mathcal{F}_{p}$ and that $\phi:(X,\cdot)\to(X,\cdot^{\prime})$ is an isomorphism. Then $\phi\mathcal{G}(X,\cdot)\phi^{-1}=\mathcal{G}(X,\cdot^{\prime})$ so in particular $\phi$ must normalize the cyclic permutation action on the blocks and thus be of the form $\phi(a,x)=(\alpha a+\beta,\pi_{a}(x))$ for some $\pi_{a}\in\operatorname{Sym}_{\mathbb{Z}_{p}}$, $\alpha\in\mathbb{Z}_{p}^{*}$, $\beta\in\mathbb{Z}_{p}$. We now calculate
@@ 4. Indecomposable irretractable cycle sets of size $p^{2}$ The general case
In this section we will now obtain a solution to the general problem:
Problem 4.1.
Let $p$ be a prime. Find all indecomposable irretractable cycle sets $X$ of size $p^{2}$.
@@
In the remainder of the section, we let $X$ be an indecomposable irretractable cycle set of size $p^{2}$ and $\mathcal{G}=\mathcal{G}(X)$. As we already covered the case where $\mathcal{G}$ is a $p$-group in Section 3, we can assume that $\mathcal{G}$ is not a $p$-group, but recall that $\mathcal{G}$ is solvable by [9, Theorem 2.14]. We associate $X$ with its image in $\mathcal{G}$, which is a transitive cycle base. As $X$ is irretractable, it follows from [2, Lemma 2.1.] that $\operatorname{Soc}(\mathcal{G})=0$.
@@
Let $q\neq p$ be prime and $\mathcal{G}_{q}$ the $q$-Sylow subgroup of $\mathcal{G}^{+}$ which is a left ideal by the same argument as above. $\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$ is a brace by 4.2 and $Y\cap\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$ is a sub-cycle set of $Y$. Now if $Y$ has finite multipermutation level, then $Y\cap\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$ is either empty or equal to $Y$ by [4, Theorem 5.1]. As $|Y|$ is a $p$-power, $\mathcal{G}_{q}$ fixes at least one point in $Y$ under the $\lambda$-action. This means that $Y\cap\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})=Y$ and thus $\mathcal{G}_{p}\subseteq\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$. It follows that $\mathcal{G}_{p}^{\circ}$ normalizes $\mathcal{G}_{q}^{\circ}$ but as $\mathcal{G}_{p}$ acts faithful and transitive on a set of size $p^{2}$, this implies that $\mathcal{G}_{q}=0$ and hence, $\mathcal{G}_{p^{\prime}}=0$ but this contradicts the assumption that $\mathcal{G}$ is not a $p$-group. We therefore deduce that $Y$ is not of finite multipermutation level. Together with the earlier observation that $|Y|\in\{1,p,p^{2}\}$, we conclude that $Y$ is irretractable of size $p^{2}$ and therefore as described in 3.8. In particular, we find that $\mathcal{G}_{p}^{\circ}$ is not abelian.
From now on we consider the unique block system of $X$ under the action of $\mathcal{G}$, and recall that the uniqueness is guaranteed by 1.7. As before we denote this by $\{\mathcal{B}_{x}\mid x\in X\}$. We denote the subgroup of $\mathcal{G}^{\circ}$ that fixes the blocks setwise by $\mathcal{A}$. Also, we define $\mathcal{A}_{p}=\mathcal{A}\cap\mathcal{G}_{p}$ and $\mathcal{A}_{p^{\prime}}=\mathcal{A}\cap\mathcal{G}_{p^{\prime}}$. Note that $\mathcal{A}_{p}$ is a $p$-Sylow subgroup of $\mathcal{A}$ and $\mathcal{A}_{p^{\prime}}$ is a Hall $p^{\prime}$-subgroup of $\mathcal{A}$. In particular, $\mathcal{A}_{p}$ is normal in $\mathcal{G}^{\circ}$ by 1.8.
@@
By 3.6 we know that $\mathcal{A}_{p}$ contains a unique representative of each block in the block system of $Y$ under the action of $\mathcal{G}_{p^{\prime}}^{\circ}$, hence also under the action of $\mathcal{G}^{\circ}$. This means that $\mathcal{A}_{p^{\prime}}$ acts trivially on the set $Y\cap\mathcal{A}_{p}$ and thus $Y\cap\mathcal{A}_{p}\subseteq\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})$. From 4.2 we know that $\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})$ is a subbrace, since $\mathcal{A}_{p^{\prime}}=\mathcal{A}\cap\mathcal{G}_{p^{\prime}}$. In particular, $\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})\cap Y$ is a sub-cycle set of $Y$ which contains $Y\cap\mathcal{A}_{p}$, but from 3.9 it then follows that $\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})\cap Y=Y$. However, as the $\lambda$-action of $\mathcal{G}$ on $Y$ is faithful, this implies that $\mathcal{A}_{p^{\prime}}=0$.
By 1.8 we find that $|\mathcal{G}_{p^{\prime}}|<p$ and thus the $\lambda$-action of $\mathcal{G}_{p}$ on $\mathcal{G}_{p^{\prime}}$ is trivial. As this same action is transitive on $Z$ we find $|Z|=1$. Let $z\in Z$, then $\lambda_{z}$ is a brace automorphism of $\mathcal{G}_{p}$ by 1.1 and therefore also its restriction to $Y$ yields a cycle set automorphism. For any $x_{i},x_{j}\in X$ we find
@@
We find that the cycle set structure on $X$ is obtained by deforming the cycle set structure on $Y$ by an automorphism of $Y$. 4.5 shows that such a deformation is always possible.
@@
Let $q\neq p$ be a prime and let $\mathcal{G}(X,\cdot)_{q}$ be the $q$-Sylow subgroup of $\mathcal{G}(X,\cdot)^{+}$. If we consider the $\lambda$-action of $\mathcal{G}(X,\cdot)_{q}$ on $Y$ we find that it has fixed points. By 4.2 we find that $\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q})$ is a brace thus $\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q})\cap Y$ is a sub-cycle set. As this sub-cycle set is non-empty, [4, Theorem 5.1] implies that $Y\subseteq\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q})$ and hence $\mathcal{G}(X,\cdot)_{p}\subseteq\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q}$. As a result, $\mathcal{G}(X,\cdot)_{q}$ is a normal subgroup of $\mathcal{G}(X,\cdot)^{\circ}$, but this is impossible as $\mathcal{G}(X,\cdot)^{\circ}$ acts transitively and faithfully on a set of $p$-power order. We conclude that $\mathcal{G}(X,\cdot)$ is a $p$-group and therefore $k=1$. ∎
@@
One can verify directly that $(X,\cdot_{\phi})$ satisfies (C1)-(C3). However, as this is useful in the proof of 4.6 we construct $(X,\cdot_{\phi})$ as the sub-cycle set of a cycle set coming from a brace.
By the functoriality of $G(X,r)$ we get an induced automorphism $\phi^{\prime}$ of $G(X,r)$ which restricts to $\phi$ on the generating set $X\subseteq G(X,\cdot)$. We let $\mathbb{Z}$ denote the trivial brace on $\mathbb{Z}$ and we let $\mathbb{Z}^{\circ}$ act on $G(X,\cdot)$ by $1\cdot g=(\phi^{\prime})^{-1}(g)$ for $g\in G(X,\cdot)$. Hence we can construct the semi-direct product $G(X,\cdot)\rtimes\mathbb{Z}$ in the sense of [20, Corollary 3.37]. The set $X+1\subseteq G(X,\cdot)\rtimes\mathbb{Z}$ is closed under the $\lambda$-action. Hence it is a sub-cycle set which is precisely the cycle set in the statement under the correspondence $x\mapsto x+1$.
@@
Let $G(X,\cdot)\rtimes\mathbb{Z}$ be as described in the proof of 4.5. The universal property of $G(X,\cdot_{\phi})$ provides a brace homomorphism $f:G(X,\cdot_{\phi})\to G(X,\cdot)\rtimes\mathbb{Z}$ mapping $x\to x+1$. The canonical map $g:(G(X,\cdot)\rtimes\mathbb{Z})^{\circ}\to\operatorname{Sym}_{X}$ sending $x\mapsto\sigma^{-1}_{x}$ and $1\mapsto\phi^{-1}$ is a group homomorphism. Clearly $\ker gf=\operatorname{Soc}(G(X,\cdot_{\phi}))$ and the image is precisely $\mathcal{G}(X,\cdot_{\phi})$. As $\mathcal{G}(X,\cdot)\cap\langle\phi\rangle=\{\operatorname{id}\}$ we find that $\mathcal{G}(X,\cdot_{\phi})^{\circ}=\mathcal{G}(X,\cdot)^{\circ}\rtimes\langle\phi\rangle$. Also, as $\ker g=\operatorname{Soc}(G(X,\cdot))\times m\mathbb{Z}$, we find that $\mathcal{G}(X,\cdot_{\phi})\cong\mathcal{G}(X,\cdot)\rtimes\mathbb{Z}_{m}$ as braces.
@@
As $\mathcal{G}(X,\cdot)$ is a Hall-subgroup of $\mathcal{G}(X,\cdot_{\phi})^{+}$, the last part of the statement follows from 4.3. ∎
@@
Assume that $f:(X,\cdot_{\phi})\to(X,\cdot^{\prime}_{\psi})$ is an isomorphism. From 4.6 and the assumptions we find the existence of some $k\in\mathbb{Z}$ such that $(X,(\cdot_{\phi})_{k})=(X,\cdot)$ and $(X,(\cdot^{\prime}_{\psi})_{k})=(X,\cdot^{\prime})$. The functoriality of cabling now yields that $f$ induces an isomorphism $f:(X,\cdot)\to(X,\cdot^{\prime})$. For any $x,y\in X$ we find $f(x\cdot_{\phi}y)=f\phi(x\cdot y)$ and $f(x)\cdot_{\psi}f(y)=\psi f(x\cdot y)$, hence $f(x\cdot_{\phi}y)=f(x)\cdot_{\psi}f(y)$ if and only if $\psi=f\phi f^{-1}$. This proves one implication of the statement.
@@
If $\alpha=1$, then the cycle sets are the ones that appear in 3.10. If $\alpha\neq 1$, then any automorphism of $(X,\cdot)$ is of the form $(a,x)\mapsto(\gamma a,\gamma x)$ for some $\gamma\in\mathbb{Z}_{p}^{*}$ with ${}^{\gamma}\Phi=\Phi$.
@@
From the discussion preceding 4.5 we know that $(X,\cdot)$ can be obtained by starting from a cycle set structure on $X$ whose permutation group is a $p$-group and deforming such cycle set by an automorphism of order coprime to $p$, in the sense of 4.5. From 3.10 it follows that, up to a cycle set isomorphism, $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and
@@
for some $\Phi\in\mathcal{R}_{p}$, $\alpha\in\mathbb{Z}_{p}^{*}$ and $\beta\in\mathbb{Z}_{p}$, satisfying ${}^{\alpha}\Phi=\Phi$. By 4.7 we may even assume $\beta=0$. We therefore get that up to isomorphism the multiplication on $X$ is precisely as in the statement.
Conversely, it follows directly from 4.5 and 4.6 that $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ with the given multiplication always yields an indecomposable irretractable cycle set. As a consequence of 4.7 we find that different choices of $\alpha$ and $\Phi$ yield non-isomorphic solutions and also that the automorphisms are the ones described in the statement. ∎

### v3

@@ 3.1. Constructing the solutions
Let $0\neq f\in\operatorname{Fix}(\mathcal{G})$. Then ${}^{f}\mathcal{B}_{x}=\lambda_{f}(\mathcal{B}_{x})\neq\mathcal{B}_{x}$ as $f\not\in\mathcal{A}$. Therefore, $\operatorname{Fix}(\mathcal{G})$ acts transitively on the system of blocks by conjugation. Furthermore, by 3.5, $\operatorname{Fix}(\mathcal{G})$ acts transitively on every single block by right-multiplication. Therefore, each $y\in X$ is of the form $y={}^{f}x\circ f^{\prime}$ for some $f,f^{\prime}\in\operatorname{Fix}(\mathcal{G})$, which can be rewritten as $y=f\circ x\circ f^{\prime}\circ\overline{f}$. ∎
From 3.6 it therefore follows that we can coordinatize the elements of $X$ in the following way: by 3.5 we can choose an element $\alpha_{0,0}\in X\cap\mathcal{A}$ and some $0\neq f\in\operatorname{Fix}(\mathcal{G})$. For $(a,x)\in\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, we then set $\alpha_{a,x}=\left({}^{\overline{f}^{\circ a}}\alpha_{0,0}\right)\circ f^{\circ x}$. By 3.6, this gives a unique coordinatization of the elements in $X$.
@@
where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(A)=\Phi(-A)$, for all $A\in\mathbb{Z}_{p}$. Vice versa, this construction always results in an indecomposable, irretractable cycle set with $\mathcal{G}(X)$ a $p$-group.
@@
A comparison shows that in order for $X$ to satisfy Eq. C1, $\Phi(b-a)=\Phi(a-b)$ must hold for all $a,b\in\mathbb{Z}_{p}$ which amounts to saying that $\Phi(A)=\Phi(-A)$, for all $A\in\mathbb{Z}_{p}$. By the same calculation, one sees that this obstruction to $\Phi$ is sufficient for $X$ to satisfy Eq. C1. By construction, all maps $\sigma_{(a,x)}$ are bijective. Furthermore, the square map
@@
Note that, using the explicit form in 3.7, we are considering the set $X\cap\mathcal{A}=\{(a,0)\mid a\in\mathbb{Z}_{p}\}$. As $\Phi$ is non-constant, we easily see that these elements generate the whole cycle set $X$. ∎
@@ 3.2. Getting rid of redundancy and determining automorphisms
Let $\mathcal{F}_{p}$ be the set of all non-constant functions $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ with the property that $\Phi(A)=\Phi(-A)$, for all $A\in\mathbb{Z}_{p}$. $\mathcal{F}_{p}$ is acted upon by $\mathbb{Z}_{p}^{\ast}$ via $({}^{\alpha}\Phi)(A)=\alpha^{-1}\Phi(\alpha A)$. From now on, let $\mathcal{R}_{p}$ be a fixed system of representatives for this action.
Recall from 3.7 that the cycle sets in the considered case are described as $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ with the operation
@@
By 3.1, $\mathcal{G}(X)^{\circ}$ is non-abelian and, by 1.6, has a unique non-trivial system of imprimitivity that consists of the blocks $\{a\}\times\mathbb{Z}_{p}$. Assume that $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ comes with two cycle set operations $\cdot,\cdot^{\prime}$ that are given by the parameters $\Phi,\Phi^{\prime}\in\mathcal{F}_{p}$ and that $\phi:(X,\cdot)\to(X,\cdot^{\prime})$ is an isomorphism. Then $\phi\mathcal{G}(X,\cdot)\phi^{-1}=\mathcal{G}(X,\cdot^{\prime})$ so in particular $\phi$ must normalize the cyclic permutation action on the blocks and thus be of the form $\phi(a,x)=(\alpha a+\beta,\pi_{a}(x))$ for some $\pi_{a}\in\operatorname{Sym}_{\mathbb{Z}_{p}}$, $\alpha\in\mathbb{Z}_{p}^{*}$, $\beta\in\mathbb{Z}_{p}$. We now calculate
@@ 4. Indecomposable irretractable cycle sets of size $p^{2}$ The general case
In this section, we will focus on the general problem. More precisely, we will construct all indecomposable irretractable cycle sets $X$ of size $p^{2}$, where $p$ is a prime number.
@@
In the remainder of the section, we let $X$ be an indecomposable irretractable cycle set of size $p^{2}$ and $\mathcal{G}=\mathcal{G}(X)$. As we already covered the case where $\mathcal{G}$ is a $p$-group in Section 3, we can assume that $\mathcal{G}$ is not a $p$-group, but recall that $\mathcal{G}$ is solvable by [9, Theorem 2.15]. We associate $X$ with its image in $\mathcal{G}$, which is a transitive cycle base. As $X$ is irretractable, it follows from [2, Lemma 2.1] that $\operatorname{Soc}(\mathcal{G})=0$.
@@
Let $q\neq p$ be prime and $\mathcal{G}_{q}$ the $q$-Sylow subgroup of $\mathcal{G}^{+}$ which is a left ideal by the same argument as above. $\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$ is a brace by 4.1 and $Y\cap\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$ is a sub-cycle set of $Y$. Now if $Y$ has finite multipermutation level, then $Y\cap\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$ is either empty or equal to $Y$ by [4, Theorem 5.1]. As $|Y|$ is a $p$-power, $\mathcal{G}_{q}$ fixes at least one point in $Y$ under the $\lambda$-action. This means that $Y\cap\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})=Y$ and thus $\mathcal{G}_{p}\subseteq\operatorname{Fix}_{\mathcal{G}}(\mathcal{G}_{q})$. It follows that $\mathcal{G}_{p}^{\circ}$ normalizes $\mathcal{G}_{q}^{\circ}$ but as $\mathcal{G}_{p}$ acts faithful and transitive on a set of size $p^{2}$, this implies that $\mathcal{G}_{q}=0$ and hence, $\mathcal{G}_{p^{\prime}}=0$ but this contradicts the assumption that $\mathcal{G}$ is not a $p$-group. We therefore deduce that $Y$ is not of finite multipermutation level. Together with the earlier observation that $|Y|\in\{1,p,p^{2}\}$, we conclude that $Y$ is irretractable of size $p^{2}$ and therefore as described in 3.7. In particular, we find that $\mathcal{G}_{p}^{\circ}$ is not abelian.
From now on we consider the unique block system of $X$ under the action of $\mathcal{G}$, and recall that the uniqueness is guaranteed by 1.6. As before we denote this by $\{\mathcal{B}_{x}\mid x\in X\}$. We denote the subgroup of $\mathcal{G}^{\circ}$ that fixes the blocks setwise by $\mathcal{A}$. Also, we define $\mathcal{A}_{p}=\mathcal{A}\cap\mathcal{G}_{p}$ and $\mathcal{A}_{p^{\prime}}=\mathcal{A}\cap\mathcal{G}_{p^{\prime}}$. Note that $\mathcal{A}_{p}$ is a $p$-Sylow subgroup of $\mathcal{A}$ and $\mathcal{A}_{p^{\prime}}$ is a Hall $p^{\prime}$-subgroup of $\mathcal{A}$. In particular, $\mathcal{A}_{p}$ is normal in $\mathcal{G}^{\circ}$ by 1.7.
@@
By 3.5 we know that $\mathcal{A}_{p}$ contains a unique representative of each block in the block system of $Y$ under the action of $\mathcal{G}_{p^{\prime}}^{\circ}$, hence also under the action of $\mathcal{G}^{\circ}$. This means that $\mathcal{A}_{p^{\prime}}$ acts trivially on the set $Y\cap\mathcal{A}_{p}$ and thus $Y\cap\mathcal{A}_{p}\subseteq\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})$. From 4.1 we know that $\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})$ is a subbrace, since $\mathcal{A}_{p^{\prime}}=\mathcal{A}\cap\mathcal{G}_{p^{\prime}}$. In particular, $\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})\cap Y$ is a sub-cycle set of $Y$ which contains $Y\cap\mathcal{A}_{p}$, but from 3.8 it then follows that $\operatorname{Fix}_{\mathcal{G}}(\mathcal{A}_{p^{\prime}})\cap Y=Y$. However, as the $\lambda$-action of $\mathcal{G}$ on $Y$ is faithful, this implies that $\mathcal{A}_{p^{\prime}}=0$.
By 1.7 we find that $|\mathcal{G}_{p^{\prime}}|<p$ and thus the $\lambda$-action of $\mathcal{G}_{p}$ on $\mathcal{G}_{p^{\prime}}$ is trivial. As this same action is transitive on $Z$ we find $|Z|=1$. Let $z\in Z$, then $\lambda_{z}$ is a brace automorphism of $\mathcal{G}_{p}$ by 1.1 and therefore also its restriction to $Y$ yields a cycle set automorphism. For any $x_{i},x_{j}\in X$ we find
@@
We find that the cycle set structure on $X$ is obtained by deforming the cycle set structure on $Y$ by an automorphism of $Y$. 4.4 shows that such a deformation is always possible.
@@
Let $q\neq p$ be a prime and let $\mathcal{G}(X,\cdot)_{q}$ be the $q$-Sylow subgroup of $\mathcal{G}(X,\cdot)^{+}$. If we consider the $\lambda$-action of $\mathcal{G}(X,\cdot)_{q}$ on $Y$ we find that it has fixed points. By 4.1 we find that $\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q})$ is a brace thus $\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q})\cap Y$ is a sub-cycle set. As this sub-cycle set is non-empty, [4, Theorem 5.1] implies that $Y\subseteq\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q})$ and hence $\mathcal{G}(X,\cdot)_{p}\subseteq\operatorname{Fix}(\mathcal{G}(X,\cdot)_{q}$. As a result, $\mathcal{G}(X,\cdot)_{q}$ is a normal subgroup of $\mathcal{G}(X,\cdot)^{\circ}$, but this is impossible as $\mathcal{G}(X,\cdot)^{\circ}$ acts transitively and faithfully on a set of $p$-power order. We conclude that $\mathcal{G}(X,\cdot)$ is a $p$-group and therefore $k=1$. ∎
@@
One can verify directly that $(X,\cdot_{\phi})$ satisfies (C1)-(C3). However, as this is useful in the proof of 4.5 we construct $(X,\cdot_{\phi})$ as the sub-cycle set of a cycle set coming from a brace.
By the functoriality of $G(X,r)$ we get an induced automorphism $\phi^{\prime}$ of $G(X,r)$ which restricts to $\phi$ on the generating set $X\subseteq G(X,\cdot)$. We let $\mathbb{Z}$ denote the trivial brace on $\mathbb{Z}$ and we let $\mathbb{Z}^{\circ}$ act on $G(X,\cdot)$ by $1\cdot g=(\phi^{\prime})^{-1}(g)$ for $g\in G(X,\cdot)$. Hence we can construct the semi-direct product $G(X,\cdot)\rtimes\mathbb{Z}$ in the sense of [19, Corollary 3.37]. The set $X+1\subseteq G(X,\cdot)\rtimes\mathbb{Z}$ is closed under the $\lambda$-action. Hence it is a sub-cycle set which is precisely the cycle set in the statement under the correspondence $x\mapsto x+1$.
@@
Let $G(X,\cdot)\rtimes\mathbb{Z}$ be as described in the proof of 4.4. The universal property of $G(X,\cdot_{\phi})$ provides a brace homomorphism $f:G(X,\cdot_{\phi})\to G(X,\cdot)\rtimes\mathbb{Z}$ mapping $x\to x+1$. The canonical map $g:(G(X,\cdot)\rtimes\mathbb{Z})^{\circ}\to\operatorname{Sym}_{X}$ sending $x\mapsto\sigma^{-1}_{x}$ and $1\mapsto\phi^{-1}$ is a group homomorphism. Clearly $\ker gf=\operatorname{Soc}(G(X,\cdot_{\phi}))$ and the image is precisely $\mathcal{G}(X,\cdot_{\phi})$. As $\mathcal{G}(X,\cdot)\cap\langle\phi\rangle=\{\operatorname{id}\}$ we find that $\mathcal{G}(X,\cdot_{\phi})^{\circ}=\mathcal{G}(X,\cdot)^{\circ}\rtimes\langle\phi\rangle$. Also, as $\ker g=\operatorname{Soc}(G(X,\cdot))\times m\mathbb{Z}$, we find that $\mathcal{G}(X,\cdot_{\phi})\cong\mathcal{G}(X,\cdot)\rtimes\mathbb{Z}_{m}$ as braces.
@@
As $\mathcal{G}(X,\cdot)$ is a Hall-subgroup of $\mathcal{G}(X,\cdot_{\phi})^{+}$, the last part of the statement follows from 4.2. ∎
@@
Assume that $f:(X,\cdot_{\phi})\to(X,\cdot^{\prime}_{\psi})$ is an isomorphism. From 4.5 and the assumptions we find the existence of some $k\in\mathbb{Z}$ such that $(X,(\cdot_{\phi})_{k})=(X,\cdot)$ and $(X,(\cdot^{\prime}_{\psi})_{k})=(X,\cdot^{\prime})$. The functoriality of cabling now yields that $f$ induces an isomorphism $f:(X,\cdot)\to(X,\cdot^{\prime})$. For any $x,y\in X$ we find $f(x\cdot_{\phi}y)=f\phi(x\cdot y)$ and $f(x)\cdot_{\psi}f(y)=\psi f(x\cdot y)$, hence $f(x\cdot_{\phi}y)=f(x)\cdot_{\psi}f(y)$ if and only if $\psi=f\phi f^{-1}$. This proves one implication of the statement.
@@
If $\alpha=1$, then the cycle sets are the ones that appear in 3.9. If $\alpha\neq 1$, then any automorphism of $(X,\cdot)$ is of the form $(a,x)\mapsto(\gamma a,\gamma x)$ for some $\gamma\in\mathbb{Z}_{p}^{*}$ with ${}^{\gamma}\Phi=\Phi$.
@@
From the discussion preceding 4.4 we know that $(X,\cdot)$ can be obtained by starting from a cycle set structure on $X$ whose permutation group is a $p$-group and deforming such cycle set by an automorphism of order coprime to $p$, in the sense of 4.4. From 3.9 it follows that, up to a cycle set isomorphism, $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and
@@
for some $\Phi\in\mathcal{R}_{p}$, $\alpha\in\mathbb{Z}_{p}^{*}$ and $\beta\in\mathbb{Z}_{p}$, satisfying ${}^{\alpha}\Phi=\Phi$. By 4.6 we may even assume $\beta=0$. We therefore get that up to isomorphism the multiplication on $X$ is precisely as in the statement.
Conversely, it follows directly from 4.4 and 4.5 that $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ with the given multiplication always yields an indecomposable irretractable cycle set. As a consequence of 4.6 we find that different choices of $\alpha$ and $\Phi$ yield non-isomorphic solutions and also that the automorphisms are the ones described in the statement. ∎